OpenStudy (anonymous):

the square root of x^7

OpenStudy (hba):

Remember, \[\sqrt{a}=a^{1/2}\]

OpenStudy (anonymous):

\[x^7 = x*x*x*x*x*x*x\]

OpenStudy (hba):

Similarly, \[\sqrt{a^5}=a^{5/2}\] Now you decide \[\sqrt{a^7}=\]

OpenStudy (anonymous):

x^(7/2) And remember, this is defined only for x >= 0 because x^7 is negative for x < 0

OpenStudy (anonymous):

or\[\sqrt{x*x*x*x*x*x*x} = x^3 \sqrt{x}\]

OpenStudy (anonymous):

i still don not understand

OpenStudy (hba):

Ok let us see this